mhchen:

How do I correct my mistake:

mhchen:

\[x = \sqrt{x + 2}\] \[x^2 = x+2\] \[x^2 - x - 2 = 0\] \[(x+1)(x-2) = 0\] \[x = -1 ,2\]

mhchen:

This isn't solving the equation, this is a mathematical proof. Ex: Since x=-1 would make x = sqrt(x+2) false, then this proof is a failure

TheSmartOne:

Not sure what mistake you're referring to Since we have a square root, we need to plug in our answer choices to make sure that it is indeed a solution. And sqrt(1) \(\neq\) -1 so we know that x \(\neq -1\) and that the only solution is x = 2

mhchen:

Yes but the problem is, I can't take out x=-1

mhchen:

It's a proof logic: If A then B If A and B then C or If and only if A then B B is false Then A has to be false

dude:

But rational functions have 2 solutions

mhchen:

\[x = \sqrt{x + 2}\] and \[x^2 = x+2\] is not bi conditional basically, I got to figure out how to make it biconditional

TheSmartOne:

Yeah dude, but there's an extraneous solution

dude:

\(-1=\pm \sqrt{-1+2}\)

dude:

Took me a while to type this On mobile ;(

mhchen:

I can't change the original equation, I need to change a step in the proof instead

mhchen:

This is propositional logic, not solving an equation. I'm trying to prove something not solve it.

TheSmartOne:

I see, mhm maybe @Vocaloid can give some insight

mhchen:

Update: I got the answer. It's a paragraph btw so I wont bother typing it out